OpenStudy (anonymous):

/Users/doreltibi/Desktop/Screen Shot 2012-12-31 at 11.02.38 AM.png

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Got it?

geerky42 (geerky42):



geerky42 (geerky42):

Or you can divide kite into four triangles then find the area of each of them then add them together.

OpenStudy (anonymous):

so for my problem it would be the following: 1/2 (15+7) (4+x) 1/2 (88+22x) 1/2 (110x) Area= 55x

OpenStudy (anonymous):

?

geerky42 (geerky42):

You got it wrong, x = 7, y = 15, z = 4

OpenStudy (anonymous):

so the area would be 13

geerky42 (geerky42):

133* Yeah.

OpenStudy (anonymous):

1/2 (7+7) (15+4) 1/2(14) (19) Area=1/2(266) Area = 133